The inventory records of Kuffel Co. reflected the following information for the year ended December 31, 2016:

   

Date Transaction Number
of Units
Unit
Cost
Total
Cost
1/1 Beginning inventory 150 $ 32 $ 4,800
1/24 Purchase 70 33 2,310
2/22 Sale (100 ) - -
3/7 Purchase 90 35 3,150
4/10 Purchase 140 38 5,320
6/11 Sale (100 ) - -
9/28 Purchase 50 38 1,900
12/4 Sale (100 ) - -

Required:

a. Assume that Kuffel Co. uses a periodic inventory system. Calculate cost of goods sold and ending inventory under FIFO and LIFO.

b. Assume that Kuffel Co. uses a perpetual inventory system. Calculate cost of goods sold and ending inventory under FIFO and LIFO.
Answer #1
Unit Unit Cost Total Cost
Beginning inventory 150 32 4800
Purchase 1 70 33 2310
Purchase 2 90 35 3150
Purchase 3 140 38 5320
Purchase 4 50 38 1900
Total 500 17480

a) Calculate cost of goods sold and ending inventory : Perpetual method

FIFO LIFO
Cost of goods sold 17480-7570 = 9910 17480-6450 = 11030
Ending inventory 1900+5320+350 = 7570 4800+50*33 = 6450

b) Calculate cost of goods sold and ending inventory : Periodic method

FIFO LIFO
Cost of goods sold 17480-7570 = 9910 17480-6640 = 10840
Ending inventory 1900+5320+350 = 7570 120*32+80*35 = 6640